What are the concentration of this following solutions

1. 5mL of 10ppm Fe+2 diluted in a 100mL volumetric flask

2. 10mL of 10ppm Fe+2 diluted in a 100mL volumetric flask

3. 20mL of 10ppm Fe+2 diluted in a 100mL volumetric flask

4. 25mL of 10ppm Fe+2 diluted in a 100mL volumetric flask
Answer #1

To solve given problem we can use dilution formula.

We have dilution formula, C stock\times V stock = C dilute\times V dilute.

Where, C stock is concentration of stock solution,  V stock is a volume of stock solution, C dilute is concentration fo dilute solution and V dilute is volume of dilute solution.

1) Given

C stock = 10 ppm,

V stock = 5 ml

C dilute = ?

V dilute. = 100 ml

Placing given values in above formula , we get 10 ppm \times 5 ml = C dilute\times 100 ml

Therefore, C dilute = 10 ppm \times 5 ml / 100 ml = 0.5 ppm

ANSWER : Concentration of diluted Fe 2+ solution = 0.5 ppm

2)  Given

C stock = 10 ppm,

V stock =10 ml

C dilute = ?

V dilute. = 100 ml

Placing given values in above formula , we get 10 ppm \times 10 ml = C dilute\times 100 ml

Therefore, C dilute = 10 ppm \times 10 ml / 100 ml = 1.0 ppm

ANSWER : Concentration of diluted Fe 2+ solution = 1.0 ppm

3)

Given

C stock = 10 ppm,

V stock =20 ml

C dilute = ?

V dilute. = 100 ml

Placing given values in above formula , we get 10 ppm \times 20 ml = C dilute\times 100 ml

Therefore, C dilute = 10 ppm \times 20 ml / 100 ml = 2.0 ppm

ANSWER : Concentration of diluted Fe 2+ solution = 2.0 ppm

4)

Given

C stock = 10 ppm,

V stock =25 ml

C dilute = ?

V dilute. = 100 ml

Placing given values in above formula , we get 10 ppm \times 25 ml = C dilute\times 100 ml

Therefore, C dilute = 10 ppm \times 25 ml / 100 ml = 2.5 ppm

ANSWER : Concentration of diluted Fe 2+ solution = 2.5 ppm